Consider the following reaction: A ↔ B

Using the thermodynamic data provided below at 25 °C, determine the equilibrium constant for this reaction. R = 8.314 J/mol K .

ΔGof(A) = -2.2 kJ/mol

ΔGof(B) = -4.8 kJ/mol

My key says the correct answer is 2.9, what do you use to get there?
Answer #1

\Delta G_{reaction}=G_{B}-G_{A}

\Delta G_{reaction}=(-4.8+2.2)\: kJ/mol=-2.6\: kJ/mol

\therefore \Delta G_{reaction}=(-4.8+2.2)\: kJ/mol=-2600\: J/mol

25^{0}C=(25+273)K=298\: K

We know, the relatiobn between free energy and equilibrium constant-

\Delta G=-RTlnK_{eq}

\therefore lnK_{eq}=\frac{-\Delta G}{R\times T}

\therefore K_{eq}=exp(\frac{-\Delta G}{R\times T})

\therefore K_{eq}=exp(\frac{-(-2600\: J/mol)}{8.314\: J/(mol.K)\times 298\: K})

\therefore K_{eq}=exp(1.049)

\therefore K_{eq}=2.859

\therefore K_{eq}\approx 2.9 [ANSWER]